# Exploring Causes of Representational Similarity in Machine Learning Models

### v1

@@ 4.3 Experimental setup
We run our task and dataset splitting experiments on a variety of settings. This section describes the basic model setup and how we evaluate representational similarity, then gives details about how we implement the different splitting approaches.
@@
Model training. Table 2 provides an overview of the models used in our experiments. To understand how our different overlap approaches affect representational similarity across modalities, we train image classifiers, variational autoencoders (VAEs), and small language models. We trained all models using similar hardware setup (Single NVIDIA RTX 6000 Ada Generation GPU). For the ResNet-18 models we leverage the FFCV library for accelerated data loading. All ResNet models used cross entropy loss with label smoothing at 0.1. The VAE leveraged a singular encoder and decoder with composite generative loss. The language models are a GPT-Style transformer [25] with optimizations such as mixed precision, gradient accumulation & clipping, and early stopping, with task overlap and/or data overlap varied across runs. In the Appendix, Table 4 gives full hyperparameter details, while Table 5 lists baseline model performance.
@@
Table 2: Models and datasets used in our experiments. ${}^{*}=$ ColorShapeDigit800K  was only used to train a ResNet18 model for task overlap experiments; all other models/datasets were used for both task and dataset overlap experiments.
Models Training Data Overlap Methods Dataset Task
ResNet18 [20] CIFAR-10 [30], CIFAR-100 [30] ✓  ✓
TinyImagenet [12], ColorShapeDigit800K (ours)
VAE [26] CIFAR-10 [30] ✓  ✓
nanoGPT [25] Tiny-{Stories, Codes} [15, 40], Shakespeare [9], ✓  ✓
Tiny-{WebText, TextBooks} [41, 39], WikiText [35]
[…]
@@
Implementation of splitting methods. For both splitting methods, we experiment with 3 training paradigms. We train image classifiers on CIFAR-10, CIFAR-100, and TinyImagenet; a VAE on CIFAR-10; and small language models on Shakespeare and TinyStories. We evaluate the representational similarity of models on both in-distribution (e.g. test set of training data) and out-of-distribution (e.g. unseen in training) datasets since similarity between models do not transfer uniformly from one dataset to another [8]. For image classifiers and the VAE, we use two popular datasets (MNIST [13] and SVHN [43]) as OOD datasets. For language data with dataset overlap, we train with either Shakespeare or TinyStories and used WikiText [35] as the OOD dataset; language models with task overlap are trained with three randomly chosen datasets in Table 2. For each training dataset, we train 10 pairs of models using 10 different seeds for image data (6 for text) and evaluate the representational similarity for model pairs on both in-distribution and OOD data.
@@
For task splitting method (1), we experiment with task overlap levels ranging from 0.0 to 1.0 and evaluate models on both in-distribution and OOD data as in the dataset splitting approach. For task overlap experiments with text models, we construct overlapping datasets by combining multiple text corpora, rather than simply splitting one dataset between two models. In our case, each dataset includes up to three three corpora, meaning the models see text from different writing styles during training. The task splitting is controlled by the number of overlapping corpora across the two models’ datasets. Therefore, we use the following terms for OOD data in the text splitting experiments: common = corpora used by both models during training; OOD (one model) = corpora seen by only one model; OOD (both models) = corpora not seen by either model.

### v2

@@ 4.2 Experimental setup
We run our task and dataset splitting experiments on a variety of settings, ranging from image classifiers to large language models to diffusion models.
@@
Model training. To understand how our different overlap approaches affect representational similarity across modalities, we finetune image classifiers and LLMs, and train small language models and diffusion models from scratch. Table 2 summarizes models and training datasets used in our experiments. All ResNet/ViT (He et al., 2016; Dosovitskiy et al., 2021) models used cross entropy loss with label smoothing at 0.1. The small language models are a GPT-Style transformer (Karpathy, 2023) with optimizations such as mixed precision, gradient accumulation & clipping, and early stopping, with task overlap and/or data overlap varied across runs. The diffusion UNet (von Platen et al., 2022) is trained using the score matching loss. We LoRA finetuned (Hu et al., 2022) Llama-3.2-Instruct (Touvron et al., 2023) models on 16 million tokens. Refer to §A.1.1 for more details. Table 5 and Figure 8 list baseline model performance.
@@
Table 2: Models and datasets used in our experiments. $^{*}=$ ColorShapeDigit800K  was only used to train ResNet/ViT models for task overlap experiments, since it is an image classification task; all other models/datasets were used for both task and dataset overlap experiments. We did not have enough resources to train ColorShapeDigit800K  on Diffusion UNets, which are slow to converge.
Models Training Data
Overlap Methods
Dataset Task
ResNet/ViT CIFAR-100, TinyImagenet, ✓  ✓
ColorShapeDigit800K (ours)
UNet (diffusion) CIFAR-10 ✓  ✓
nanoGPT and Tiny-{Stories, Codes}, Shakespeare, ✓  ✓
Llama-3.2-Instruct Tiny-{WebText, TextBooks}, WikiText
Measuring representational similarity. After training, we evaluate the downstream effects of the various overlap methods by measuring the representational similarity between pairs of models trained with a specific dataset or task overlap. Let $f_{\hat{\theta}}(z)$ denote the representation of an input $z$ for all $z\in\mathbb{D}$. We use a kernel matrix $\mathbf{K}:\mathbb{D}\times\mathbb{D}\rightarrow\mathbb{R}$ of size $n\times n$ to measure similarity between representations, where $\mathbf{K}(z_{i},z_{j})=\langle f_{\hat{\theta}}(z_{i}),f_{\hat{\theta}}(z_{j})\rangle$ and $\langle\cdot,\cdot\rangle$ is the inner product. Representation similarity between two models is often evaluated via kernel matrix comparison through a technique known as Centered Kernel Alignment (CKA) (Kornblith et al., 2019). Other common metrics include nearest-neighbor scores (Klabunde et al., 2025), mutual KNN (Huh et al., 2024), and SVCCA (Raghu et al., 2017). We use implementations of these from the codebase of Huh et al. (2024), and report CKA in our main results (§5) due to its widespread in both vision and language models (Huh et al., 2024; Klabunde et al., 2023). Additional metrics are evaluated in Appendix A.2.
Splitting method implementation and evaluation. We implement dataset and task splitting differently for the different modalities we explore—see §A.1 for details. For each setting, we train pairs of models with various task/dataset overlap across 4 different seeds and evaluate the representational similarity (using CKA) for model pairs on held-out validation sets from the initial training datasets.
5 Experimental Results
Overall, we find that increasing both task and dataset overlap drives higher representation alignment between models, and that combining them produces the strongest trend.
5.1 Effect of dataset overlap on representation similarity
(a) ResNets - TinyImageNet
(b) Llama - Tiny-Codes
(c) ViTs - TinyImageNet
Figure 4: Higher dataset overlap correlates with higher representational similarity across the models we evaluate. Graphs are titled as {Model architecture}. We report CKA scores, computed on held-out validation datasets, (y axis) vs. amount of dataset overlap between models (x axis).
As described previously, we control training data overlap while holding all other variables (dataset size, task, model architecture) constant. We report CKA scores between models versus models’ training dataset overlap in Figure 4. As this figure shows, when datasets overlap more, the CKA score computed between model pairs steadily rises, even across different model types, sizes, and modalities. We fit a linear model to these result, using dataset overlap as the independent variable and CKA score as the dependent and report these results in Figure 4. This trendline further confirms our results, as the fitted regressions generally have positive slope and high $R^{2}$ values. There are some exceptions in the Llama results. As described in §4.2, we fine-tuned Llama models rather than training from scratch due to compute limitations. We hypothesize that the relatively small set of trainable LoRA parameters and the small fine-tuning dataset result in insignificant changes to the model’s representations, yielding the observed lack of trend. The consistent trend in ViT/ResNet, for which we do end-to-end training, suggests that training data overlap drives representational alignment in models trained from scratch.
5.2 Effect of task overlap on representation similarity
Next, we evaluate how our two different methods of task splitting affect representational similarity. First, we consider method (1), task splitting with increasing dataset overlap, which closely mirrors the setup of our dataset splitting experiments. Then, we analyze method (2), task splitting with constant dataset overlap, which eliminates the confounding variable of varying dataset size at each task split.
(1) Task splitting with increasing dataset overlap. When both dataset and task overlap simultaneously increase, we observe a positive trend between task overlap and model CKA, see Figure 5.
(a) ResNet - TinyImageNet
(b) Llama - Tiny-Codes
(c) ViTs - TinyImageNet
Figure 5: Higher task overlap positively correlates with higher representational similarity across the models we evaluate. Results for task splitting method (1), in which dataset and task overlap increase together. Graphs are titled as {Model architecture}. We report CKA scores, computed on held-out validation datasets, (y axis) vs. amount of task overlap between models (x axis).
(a) ViTs
(b) ResNets
Figure 6: Increasing task overlap with constant data overlap positively correlates with representational similarity. Results for task splitting method (2). We report CKA scores for ViT and ResNet models trained on different overlap levels of ColorShapeDigit800K, as described in §4.2.
(2) Task splitting with constant data overlap. The first task splitting approach allows us to evaluate the same datasets/models as in the dataset splitting experiments but introduces a confounding variable since dataset overlap increases as task overlap increases. Therefore, we isolate the effect of scaling task overlap through experiments on ColorShapeDigit800K, as described in §4.2. Results from these experiments are shown in Figure 6. We find that, even when images in $D_{1}$ and $D_{2}$ are identical, increased task overlap positively correlates with increased CKA similarity scores in the models we evaluate. Linear regression models fit to these data (independent variable is task overlap, dependent variable is CKA) have a strong positive slope and reasonable $R^{2}$ values. These results empirically show that shared tasks among models, independent of data overlap, drive representational alignment.
Finally, we evaluate which of our overlap methods—dataset alone, task alone, or dataset and task together—most strongly influences downstream representational similarity. To measure this, we use mutual information (MI) (Ross, 2014), an information-theoretic measure that quantifies the amount of shared information between two random variables. We compute the weighted mean mutual information for each overlap setting, where each model architecture evaluated contributes equally. Results are shown in Table˜3. In this table, “task overlap” refers to experiments with ColorShapeDigit800K in which we vary task while holding dataset constant, while “task + dataset overlap” refers to experiments in which task and dataset overlap increase together. As Table˜3 shows, the dataset and task overlap combined method has the highest MI score. This indicates that datasets with many shared characteristics (e.g. shared data points and task objectives) yield models with the most similar feature representations.
Table 3: Mutual information between dataset/task overlap and representational similarity is highest when both dataset and task overlap. We report average mutual information between overlap and CKA metric (we use CKNNA for diffusion UNet) across overlap types, weighted by model architecture. “Task overlap” refers to task splitting method (2) with ColorShapeDigit800K, while “task + dataset overlap” refers to task splitting method (1).
Dataset Overlap Task Overlap Task + Dataset Overlap
Mutual Information 0.584 0.771 0.891
6 Downstream Consequences of Representational Similarity
Finally, we explore downstream effects from dataset/task overlap, and consider whether one overlap method more strongly influences behavioral similarity. Here, we consider two downstream behaviors: adversarial example transferability and jailbreak transferability. This exploration is motivated by prior work highlighting the possibility of transferable adversarial examples and jailbreaks (Zou et al., 2023; Demontis et al., 2019; Carlini & Wagner, 2017). While intriguing, these works do not investigate how (or if) varying levels of model representational similarity affect transfer attack success rates.
Transferable Adversarial Examples. We use the MI-FGSM attack (Dong et al., 2018) to perform transfer attacks between ResNet and ViT models using TinyImageNet for task splitting method (1) and dataset splitting. For each pair of models trained using the same seed and overlap, we generate adversarial images using the MI-FGSM attack with whitebox access to one model and evaluate the transferable attack success rate (ASR) by computing if this example fools the other model in the pair. As Figure 7 (b) shows, we observe a strong positive trend between dataset/task overlap, providing empirical evidence that high training set overlap between models may increase their susceptibility to transfer attacks. See A.3 for attack settings, whitebox ASR, example images, and full results.
Transferable Jailbreaks. To implement jailbreak attacks, we follow the methodology of Andriushchenko et al. (2025), which uses a random search algorithm to construct an adaptive jailbreak attack for arbitrary LLMs by finding an adversarial suffix that can be appended to prompt $P$ to elicit unsafe responses. We use the input prompt template and random search algorithm from Andriushchenko et al. (2025) to generate jailbreak prompts for one model in a pair, then test the jailbreak on the other model. We evaluate jailbreak attack success using the LLM-as-a-judge technique from (Chao et al., 2024; 2025), in which a secondary LLM evaluates if the jailbreak was successful. We report transferable jailbreak attack success rates (ASR) on pairs of Llama models fine-tuned with various dataset and task overlap percentages and use Llama-70B as the judge. Specifically, let $LLM_{split0}$ and $LLM_{split1}$ be models fine-tuned on the splits described in section 4.1. We evaluate cross-split attack transferability by testing how adversarial suffixes optimized for models trained on one split perform when applied to models trained on the other.
As Figure 7(a) shows, transferable jailbreak ASR remains high across various overlap percentages. We hypothesize this is because the large LLMs were fine-tuned and not trained from scratch, which resulted in smaller changes in similarity scores between models across various overlap percentages as shown in Figure 4(b). Refer to A.4 for full jailbreaking result.
Summary. Our exploration here highlights how representation similarity can be a useful indicators of security risks between a pair of machine learning models. We provide experimental evidence that high similarity scores correspond to higher transferability of both adversarial examples and jailbreak attacks, which to the best of our knowledge has not been explored in the literature.
(a) Llama 3.2, Llama-70B judge
(b) ViT
Figure 7: Transfer attack success rate (ASR) for LLMs and ViTs of various sizes. Results for task splitting method (1). We report transfer ASR of adversarial samples (prompt for LLM, perturbed image for ViTs) between pairs of models with the same overlap.
7 Discussion
Broader impacts. Today’s large-scale AI models, particularly generative models, are often trained on overlapping datasets (§3.1) for similar purposes, e.g. serving as general purpose chatbots. Our findings that dataset and task overlap drive model similarity then suggest these models may have unexpectedly similar feature space representations. Recent work has surfaced concerns about how feature space alignment between models could result in homogeneous model behaviors. For example, Wenger & Kenett (2025) showed that a broad set of generative AI models returns a narrow set of responses—much narrower than human responses—when responding to creative prompts. Homogeneity in AI models, in creative domains and beyond, could negatively impact AI users, who may find their AI-influenced behaviors unexpectedly narrowed and/or driven towards those of other AI use (Peterson, 2025).
Limitations and future work. Our work has several limitations. First, we primarily use CKA to measure representational similarity. Prior work has pointed out downsides of CKA relative to other metrics (Bansal et al., 2021; Sucholutsky et al., 2024). CKA scores are also difficult to interpret—they have maximum theoretical value of 1, but there is no established threshold for what score indicates strong alignment between models. Second, the size of models with which we experiment is limited due to compute constraints. This prevents us from making statements about how dataset and task overlap might affect similarity of today’s production-scale AI models. Third, we only compare models with the same architecture to avoid introducing confounding variables in our analysis, but future work could design controlled experiments to address this. Finally, we consider two possible factors affecting model similarity, but other factors could exist, leaving much room for future exploration.
Appendix A Appendix
A.1 Details on task/dataset splitting method implementations
For both splitting methods, we experiment with 3 training paradigms. We train and finetune image classifiers (ResNet, ViT architectures) on CIFAR-100, and TinyImagenet; train diffusion UNets using CIFAR-10; and train small language models using the nanoGPT architecture, and finetune Llama models. We evaluate the representational similarity of models the validation set of datasets.
@@ A.1 Details on task/dataset splitting method implementations
For language data with dataset overlap, we train nanoGPT models with either Shakespeare or TinyStories and we finetune Llama models with TinyCodes; language models with task overlap are trained with three randomly chosen datasets in Table˜2. For each setting, we finetune pairs of models using a few different seeds for image data and evaluate the representational similarity for model pairs on validation data.
For both task and dataset splitting across all modalities, we always use two datasets of equal size regardless of the value of $\alpha$. Under task splitting, entire partitions (in image datasets, a partition is a class; in text datasets, a partition is a contiguous text from one corpus) are assigned to one dataset or the other. For $\alpha>0$, this approach means some partitions from the original training distribution $\mathbb{D}$ are excluded from both datasets ($D_{1}$ and $D_{2}$).
@@
For task splitting method (1), we experiment with task overlap levels ranging from 0.0 to 1.0 and evaluate models on the validation set. For task overlap experiments with text models, we construct overlapping datasets by combining multiple text corpora, rather than simply splitting one dataset between two models. In our case, each dataset includes three three corpora, meaning the models see text from different writing styles during training. The task splitting is controlled by the number of overlapping corpora across the two models’ datasets.